# DeltaZip: Efficient Serving of Multiple Full-Model-Tuned LLMs

### v2

@@ 6.2. Post-Compression Model Quality
We first study how inference accuracy is impacted by applying quantization and sparsification on model deltas compared to directly on FMT model weights. As shown in Table 1, DeltaZip achieves up to 13×13$\times$13 × compression ratio with comparable accuracy to the original FP16 FMT models. On smaller models, DeltaZip achieves 8×8$\times$8 × to $\times1$ compression ratio with comparable accuracy. In contrast, SparseGPT — which applies similar techniques directly to the FMT model weights — has a substantial drop in accuracy for all models. DeltaZip also achieves comparable accuracy compared with the state-of-the-art quantization algorithm AWQ (AWQ, ) but with a much higher compression ratio. We also observe that in some cases, the accuracy of the compressed model is even higher than the full-precision model. This effect is also noticeable in other works (han2015deep, ; AWQ, ; frantar_gptq_2023, ) and is likely due to the regularizaiton effects of compression. These results show that $\Delta$Compress can aggressively compress model deltas while retaining higher accuracy than the traditional approach of compressing model weights directly. In §6.4, we also show that compressed delta serving achieves higher accuracy on complex tasks than LoRA serving.
@@ 7. Related Work
Post-Training Compression. Another line of work focuses on reducing the memory footprint of large models by compressing the model weights in a lossy manner. Beyond the OBS framework (hassibi_optimal_1993, ) and its improvements, such as GPTQ (frantar_gptq_2023, ) and SparseGPT (frantar_sparsegpt_2023, ), many compression algorithms (AWQ, ; xiao_smoothquant_2023, ; yao_zeroquant_2022, ; tseng_quip_2024, ) have been proposed to reduce the model size. Such techniques can also be applied to model deltas by adapting the DeltaZip compression pipeline. There is also concurrent work on model delta compression. DeltaZip (yao2023deltazip, ) and GPT-Zip (isik_gpt-zip_2023, ) proposes quantization and unstructured sparsity, BitDelta (liu2024bitdelta, ) proposes extreme quantization and DARE (yu2024language, ) proposes unstructured sparsity. Compared to these works, DeltaZip optimizes more for hardware-efficiency by combining quantization with structured sparsity. To the best of our knowledge, DeltaZip is also the first serving system to support both LoRA and compressed delta FMT model serving.
@@ 8. Discussion
Limitations. While DeltaZip’s decoupled computation improves throughput and reduces latency when concurrently serving many models, decoupled inference still has higher unloaded latency than serving a FMT model directly in GPU memory. Hence, when there are only a handful of models to serve and they fit in GPU memory, DeltaZip may not be suitable. In addition, the co-serving of LoRA and FMT models in DeltaZip is at a coarse granularity, where LoRA and FMT models are served in two separate sets of GPUs and must be in separate batches. We plan to explore the possibility of serving LoRA and FMT models in the same batch as future work.
@@ 9. Conclusion
To conclude, DeltaZip efficiently serves a variety of fine-tuned model variants, whether they are fine-tuned through parameter-efficient or full-model tuning techniques. For efficient serving of full-model-tuned models, DeltaZip leverages a key insight: fine-tuning typically results in small perturbations, allowing model deltas to be highly compressible. DeltaZip co-designs the serving system with the compression algorithm and achieves 10$\times$ compression ratio, improves serving throughput by 2$\times$ to 12$\times$ and maintains high model quality comparable to FP16 models. Our system will be open-source upon publication.
@@ Acknowledgements.
We would like to thank Berivan Isik, Hermann Kumbong, Wanyi Ning, Ce Zhang, Sanmi Koyejo for their contribution and insightful feedback during the early stage of the project.

### v3

@@ 6.2. Post-Compression Model Quality
We first study how inference accuracy is impacted by applying quantization and sparsification on model deltas compared to directly on FMT model weights. As shown in Table 1, DeltaZip achieves up to 13×13$\times$13 × compression ratio with comparable accuracy to the original FP16 FMT models. On smaller models, DeltaZip achieves 4×4$\times$4 × to $\times1$ compression ratio with comparable accuracy. We also found the compression ratio for Gemma 2 models is lower than the Llama models. This is because the Gemma 2 models have a relatively larger number of parameters stored in the embedding layers, which are not compressed. In contrast, SparseGPT — which applies similar techniques directly to the FMT model weights — has a substantial drop in accuracy for all models. DeltaZip also achieves comparable accuracy compared with the state-of-the-art quantization algorithm AWQ (AWQ, ) but with a much higher compression ratio. We also observe that in some cases, the accuracy of the compressed model is even higher than the full-precision model. This effect is also noticeable in other works (han2015deep, ; AWQ, ; frantar_gptq_2023, ) and is likely due to the regularization effects of compression. These results show that $\Delta$Compress can aggressively compress model deltas while retaining higher accuracy than the traditional approach of compressing model weights directly. In §6.4, we also show that compressed delta serving achieves higher accuracy on complex tasks than LoRA serving.
@@ 7. Related Work
Post-Training Compression. Another line of work focuses on reducing the memory footprint of large models by compressing the model weights in a lossy manner. Beyond the OBS framework (hassibi_optimal_1993, ) and its improvements, such as GPTQ (frantar_gptq_2023, ) and SparseGPT (frantar_sparsegpt_2023, ), many compression algorithms (AWQ, ; xiao_smoothquant_2023, ; yao_zeroquant_2022, ; tseng_quip_2024, ) have been proposed to reduce the model size. Such techniques can also be applied to model deltas by adapting the DeltaZip compression pipeline. There is also concurrent work on model delta compression. The initial version of our system (yao2023deltazip, ) and GPT-Zip (isik_gpt-zip_2023, ) proposes quantization and unstructured sparsity, but as this paper shows, with structured sparsity, we can achieve similar compression ratio and further improve the system performance. BitDelta (liu2024bitdelta, ) proposes extreme quantization and DARE (yu2024language, ) proposes unstructured sparsity. Compared to these works, DeltaZip optimizes more for hardware efficiency by combining quantization with structured sparsity. To the best of our knowledge, DeltaZip is also the first serving system to support both LoRA and compressed delta FMT model serving.
@@ 8. Discussion
Limitations. While DeltaZip’s decoupled computation improves throughput and reduces latency when concurrently serving many models, decoupled inference still has higher unloaded latency than serving a FMT model directly in GPU memory. Hence, when there are only a handful of models to serve and they fit in GPU memory, DeltaZip may not be suitable. Compared to dedicated instances for each model, DeltaZip may be less performant, but it is more cost-effective and are more suited for environments where the cost and performance needs to be balanced. In addition, the co-serving of LoRA and FMT models in DeltaZip is at a coarse granularity, where LoRA and FMT models are served in two separate sets of GPUs and must be in separate batches. We plan to explore the possibility of serving LoRA and FMT models in the same batch as future work. Furthermore, DeltaZip reorders requests to to group requests belonging to the same delta together, thus it cannot guarantee the SLO constraints of individual models. Potential future work includes adding mechanisms to prioritize models based on their constraints, as well as predicting the output length of each request to better guarantee SLOs. Finally, the starvation handling mechanism in DeltaZip is simple and may not be beneficial for all cases. For example, the preemption of requests that are about to finish leads to unnecessary starvation and performance degradation. We plan to explore more sophisticated mechanisms, such as output length prediction, to handle starvation in future work.
@@ 9. Conclusion
To conclude, we propose DeltaZip, a serving system that enables LLM service providers to efficiently serve multiple fine-tuned models, whether they are fine-tuned through parameter-efficient or full-model tuning techniques. For efficient serving of full-model-tuned models, DeltaZip leverages a key insight: fine-tuning typically results in small perturbations, allowing model deltas to be highly compressible. DeltaZip co-designs the serving system with the compression algorithm and achieves 10$\times$ compression ratio, improves serving throughput by 2$\times$ to 12$\times$ and maintains high model quality comparable to FP16 models. One practical use case of DeltaZip is to pack less-popular models on a limited pool of GPUs, and for those popular models, service providers can still deploy them on dedicated machines, which allows for better resource allocation and minimizes operational costs.
@@ Acknowledgements.
We would like to thank Berivan Isik, Hermann Kumbong, Wanyi Ning, Ce Zhang, Sanmi Koyejo for their contribution and feedback during the early stage of the project. We also thank Lijie Xu, Yongjun He, Wenqi Jiang, Foteini Strati, anonymous EuroSys reviewers, and our shepherd, Thaleia Dimitra Doudali, for their insightful reviews and feedback. This work is partially supported by a Google Academic Research Award.
Appendix A Artifact Appendix
A.1. Abstract
A.2. Description & Requirements
Our open-source code for artifact evaluation can be found at: https://github.com/eth-easl/deltazip-ae as well as https://doi.org/10.5281/zenodo.14870277. We also provide a docker image (ghcr.io/xiaozheyao/deltazip) for easy setup of the software environment. The README file in the repository provides instructions on how to use the code and produce results.
A.2.1. How to access
Our code is available at: https://github.com/eth-easl/deltazip-ae. The docker image is available at: ghcr.io/xiaozheyao/deltazip.
A.2.2. Hardware dependencies
For the compression experiments, our reproduce script requires a machine with NVIDIA ampere GPUs (e.g., A100, RTX 3090, etc.) with equal or more than 24GB GPU memory. The machine should have at least 100 GB of storage to store the model weights and compressed models.
For the performance experiments, our reproduce script requires a machine with 4 NVIDIA A100 GPUs, and over 1TB of storage. The machine should have at least 480GB of RAM.
A.2.3. Software dependencies
We provide a docker image for easy setup of the software environment. In order to run the docker image, the host machine should have Docker and nvidia-container-toolkit installed. The host machine should also have installed CUDA 12.0 and above.
If you prefer to run the code without the docker image, the software dependencies can be found in the requirements.txt file under compression and serving directories. We recommend refer to the Dockerfile for the exact dependencies of our artifact.
Notably, we also rely on lm-eval-harness to evaluate the accuracy of the models.
A.2.4. Benchmarks
A.3. Set-up
Please set up the environment variables and paths using scripts/env.sh.
A.4. Evaluation workflow
Our evaluation workflow consists of two parts: compression/model quality evaluation and performance evaluation.
A.4.1. Major Claims
Our major claims are that the model deltas can be compressed by a factor of 10x into 2-bit per parameter and 50% structured sparsity, with minimal loss in model quality. Such compression can be leveraged to improve the performance of the model serving system.
A.4.2. Experiments
Experiment (E1): [Compressing 7B Model] [30 human-minutes + 3 compute-hours]: We compress the 7B model to 2-bit per parameter and 50% structured sparsity. We expect the compressed model to have a compression ratio of 10x with minimal loss in model quality.
    bash scripts/1_compress_7b.sh
The above command should take around 30 minutes to run.
We then test that the model can generate reasonable text samples.
    bash scripts/1.1_test_generate.sh
We then evalaute the model quality using the lm-eval-harness toolkit, by running the following command:
    bash scripts/2_eval_quality_compressed.sh
The above commands will print the accuracy of the model in a table format.
To compare the model quality of the compressed model with the original model, we evaluate the performance of the original model using the following command:
    bash scripts/2_eval_quality_original.sh
To compare the model quality with other baseline compression techniques, we first compress the models using SparseGPT and AWQ, by running the following commands:
    bash scripts/2.1_compress_baselines.sh
Then we can evaluate the model quality of the compressed models using the following command:
    bash scripts/2.2_eval_baselines.sh
We provide a helper script to summarize the results:
python scripts/helpers/aggregate_accuracy.py --compressed-model \$WORKDIR/compressed_models/lmsys.vicuna-7b-v1.5.2b_2n4m_128bs --full-model lmsys/vicuna-7b-v1.5 --accuracy-dir \$WORKDIR/eval_results --sparsegpt-model \$WORKDIR/sparsegpt_models/lmsys.vicuna-7b-v1.5.4b_2n4m_128bs --awq-model \$WORKDIR/awq_models/awq.lmsys.vicuna-7b-v1.5.4b128g
Experiment (E2): [Performance Evaluation] [60 human-minutes + 5 compute-hours]:
In order to evaluate the serving performance of our system, we first need to compress the 13b model, prepare the delta weights and create multiple copies.
    bash scripts/3_compress_13b.sh
    bash scripts/3_prepare_delta.sh
We pre-partition the model weights into the tensor-parallel shards. If you change the tensor-parallel degree, you need to re-partition the model weights. By default, we set tensor parallel degree to be $4$.
In another terminal, run the following command to start the client:
python3 scripts/helpers/bench.py --workload scripts/workload/azure.ar=0.5.jsonl --base-model meta-llama/Llama-2-13b-hf --output \$WORKDIR/results
Note that the performance numbers may vary depending on the exact machine you are running the evaluation on. In general, we expect DeltaZip to have lower latency and higher throughput compared to the baseline.